# Evaluating AI-Generated Essays with GRE Analytical Writing Assessment

### v2

@@ Abstract
The recent revolutionary advance in generative AI enables the generation of realistic and coherent texts by large language models (LLMs). Despite many existing evaluation metrics on the quality of the generated texts, there is still a lack of rigorous assessment of how well LLMs perform in complex and demanding writing assessments. This study examines essays generated by ten leading LLMs for the analytical writing assessment of the Graduate Record Exam (GRE). We assessed these essays using both human raters and the e-rater® automated scoring engine as used in the GRE scoring pipeline. Notably, the top-performing Gemini and GPT-4o received an average score of 4.78 and 4.67, respectively, falling between ”generally thoughtful, well-developed analysis of the issue and conveys meaning clearly” and ”presents a competent analysis of the issue and conveys meaning with acceptable clarity” according to the GRE scoring guideline. We also evaluated the detection accuracy of these essays, with detectors trained on essays generated by the same and different LLMs.
Keywords: Large Language Models, Writing Assessment, GRE
@@ 1 Introduction
Recent research (W.-L. Chiang et al., 2023; OpenAI et al., 2024; Touvron, Lavril, et al., 2023; Touvron, Martin, et al., 2023) has shown that extensive pretraining on diverse corpora of human-authored texts and subsequent refinement via human feedback (Ouyang et al., 2022) have enabled large language models (LLMs) to produce coherent and contextually relevant writing products. Despite these significant advancements, AI-generated texts are not necessarily perfect. Several ongoing community efforts, such as the General Language Understanding Evaluation (GLUE) (A. Wang, 2018), its enhanced version SuperGLUE (A. Wang et al., 2019), and the more recent Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), are evaluating the quality of LLM-generated texts to benchmark the progress. However, most of the aforementioned benchmarks are designed in multi-choice Question Answering format, and LLMs optimized for benchmark performances are less robust against simple dataset perturbation Alzahrani et al. (2024). Meanwhile, there are also efforts to use LLMs to evaluate LLM-generated texts automatically. For example, Fu et al. (2023) proposed an LLM-based evaluation protocol that interprets the model-predicted probability of specific token(s) as a quality score, while C.-H. Chiang & Lee (2023) and Liu et al. (2023) prompted LLMs to produce a Likert-style score ranging from 1 to 5 to assess LLM-generated texts.
From an assessment perspective, most of these evaluation metrics are based on relatively simple tasks, such as question answering, summarization, next-word prediction, and so on. There are also findings that suggest the LLM-based automated metrics could be unfaithful (P. Wang et al., 2023; Shen et al., 2023). As such, there is still a lack of rigorous evaluation of AI-generated writing products based on challenging writing tasks with a well-established scoring procedure. This study addresses this gap by leveraging the Graduate Record Examination (GRE) analytical writing assessment to evaluate AI-generated essays. The GRE analytical writing is well known for assessing the test-taker’s ability to think critically and communicate complex ideas effectively. It employs well-trained human raters and automated scoring engines, e-rater® (Attali & Burstein, 2006), to appraise written responses based on a multifaceted scoring framework (Educational Testing Service, n.d.-a). The assessment aims to gauge a candidate’s prowess in articulating and substantiating complex thoughts, constructing and assessing arguments, and maintaining a coherent and focused discourse throughout the essay.
We instructed several state-of-the-art LLMs to produce essays in response to GRE writing prompts. The selection of LLMs for this study encompasses the latest in AI technology as of mid-2024, including GPT-4o, Gemini, and Llama3-8b, as well as prior models dated back to mid-2023, including GPT-4, GPT-3.5 turbo, Google’s Bard, Llama, Vicuna, and Koala. We included these earlier models to explore potential trends that might emerge over time. All the generated essays were then evaluated by human raters and ETS’s automated scoring engine, e-rater®. By comparing the scores and language feature metrics of the generated essays, we aim to gain insights into AI-generated essays through a rigorous scoring framework set by the GRE program.
On the other hand, as AI technology advances, it becomes increasingly challenging to distinguish between human-written and AI-generated content (Tulchinskii et al., 2023; Weber-Wulff et al., 2023; Wu et al., 2023). This presents significant concerns, particularly in academic integrity and the preservation of authentic authorship. The detection accuracy of AI-generated texts is influenced by both the context and the contrast samples used. For large-scale writing assessments, AI-generated essays have been shown to be detectable with high accuracy (Hao & Fauss, 2024; Hao et al., 2024; Yan et al., 2023; Jiang et al., 2024a, b). However, as the number of high-performing LLMs continues to grow, it becomes increasingly important to benchmark the detectability of AI-generated essays across different LLMs and to assess the performance of detectors trained on various datasets.
This study aims to answer the following three research questions.
  • RQ1. What is the score distribution when LLM-generated essays are evaluated using the GRE analytical writing scoring framework?
  • RQ2. How do essays generated by different LLMs compare in terms of adherence to prompt-specified length, language features, textual similarity, and perplexity?
@@
  • RQ3. How effective are the detectors in identifying LLM-generated essays using language and perplexity features when trained on essays produced by the same or different LLMs as those in the evaluation dataset?
@@ 2 Methods
2.1 Selection of LLMs
This study aims to evaluate the quality of essays generated by a spectrum of advanced LLMs, encompassing proprietary and open-source technologies, by May 20, 2024. Within the proprietary sector, we chose the top ones, including GPT-4o from OpenAI and Gemini-1.5 from Google. These closed-source models represent the forefront of AI research in the industry, offering a glimpse into the state-of-the-art language generation capabilities. In addition to these new models, we also included some older ones, such as GPT-4 and GPT-3.5 turbo from OpenAI and Bard from Google, which appeared in the middle of 2023.
On the other hand, we also included some leading open-source LLMs in our evaluation. Specifically, we considered the recent Llama-3 8B model, which has been the subject of extensive research since its release (Dubey et al., 2024). In addition, we experimented with some older models from the middle of 2023, including the Llama-1 13B (Touvron, Lavril, et al., 2023), as well as its two finetuned variants: Vicuna (W.-L. Chiang et al., 2023) and Koala (Geng et al., 2023), both of which underwent specialized optimization processes to further refine their language generation mechanisms. The 13-billion parameter version was selected for each of these earlier models. In addition to the Llama family of models from Meta, we included another popular LLM, Mistral-7B-Instruct-v0.1, from Mistral AI, which is comparable in size to the Llama-3 model. We include model details in Appendix B.
2.2 Prompts and Essay Generation
The GRE analytical writing consists of a 30-minute writing task, ”analyze an issue”, which requires examinees to ponder a topic, evaluate its intricacies, and formulate a well-reasoned argument buttressed by examples and explanations. All writing prompts of the GRE writing assessment are publicly available (ETS, n.d.-b). To alleviate the risk of biases introduced by an individual writing prompt, we randomly selected two writing prompts for this study. To prompt LLMs for generating the essays, we sent the original GRE writing prompt plus an instruction on length: ”Please keep the response to about 500 words.”, where the 500 words is the typical length of GRE essays. The resulting prompts are listed below.
We generated 100 essays for each prompt from each LLM, leading to a total of 2,000 essays from the ten models. Each LLM comes with a set of hyperparameters that can be adjusted to control the text generation. Among these hyperparameters, temperature directly controls the variability of the generated texts. For comparison purposes, we set the temperature to 0.4 (This was used as the default value in earlier versions of the web-based ChatGPT to strike a balance between diversity and quality) for each model except Bard, where this parameter configuration was not accessible. Here, we caution that different LLMs may use varying scaling for their temperature, meaning that the same numerical value may affect the randomness and diversity of responses differently across models.
2.3 Evaluation and Detection
To address the first research question, we evaluated the essays based on the GRE scoring framework. The GRE program has a well-established scoring process involving well-trained human raters and an automated scoring engine, e-rater® (Educational Testing Service, n.d.-a). The human raters in this study were professional scoring experts in the GRE program, and they rated each essay with a score from 0 to 6 (see Appendix for more details). The e-rater® engine takes in an essay and outputs a final score as well as a set of language feature scores.
To address the second research question, we first checked how well the length of the generated essays adhered to our instructions in the prompt, as essay length is often positively correlated with essay scores in writing assessments. Then, we compared the pairwise similarity (semantic and verbatim) of generated essays, reflecting the diversity of the texts generated by the given LLMs. We then compared the e-rater® computed language features of these generated essays, such as grammar, usage, mechanics, styles, organization, development, and word complexity Attali & Burstein (2006); J. Chen et al. (2016). Table 1 below lists the definitions of a subset of the e-rater® language features we used in this study. Finally, we compared the perplexity of the generated essays, which characterizes how likely a sequence of words (essay) is generated from a given language model. Most LLMs used in this study do not allow the direct computation of perplexity. So, we chose GPT2 as a common language model to compute the perplexity of essays generated from all LLMs. As such, the perplexity here is a relative measure of how likely the essays have been generated by GPT2. Prior research has shown that perplexity based on a common language model is very informative for detecting AI-generated essays (e.g., Hao & Fauss (2024); Jiang et al. (2024a)).
To address the third research question, we randomly sampled 100 human-written essays corresponding to each of the above two prompts from the GRE writing test in 2021 when ChatGPT-like AIs were not available. These samples were mixed with essays generated by the LLMs discussed earlier. Based on the e-rater® engine-generated language features and perplexity-based features from these data, we developed machine learning classifiers to detect AI-generated essays and evaluate their performance.

### v4

# AI-generated Essays: Characteristics and Implications on Automated Scoring and Academic Integrity

@@ Abstract
The rapid advancement of large language models (LLMs) has enabled the generation of coherent essays, making AI-assisted writing increasingly common in educational and professional settings. Using large-scale empirical data, we examine and benchmark the characteristics and quality of essays generated by popular LLMs and discuss their implications for two key components of writing assessments: automated scoring and academic integrity. Our findings highlight limitations in existing automated scoring systems, such as e-rater®, when applied to essays generated or heavily influenced by AI, and identify areas for improvement, including the development of new features to capture deeper thinking and recalibrating feature weights. Despite growing concerns that the increasing variety of LLMs may undermine the feasibility of detecting AI-generated essays, our results show that detectors trained on essays generated from one model can often identify texts from others with high accuracy, suggesting that effective detection could remain manageable in practice.
Keywords: Automated Scoring, AI-generated Essay, Detection
@@ 1 Introduction
Writing is a foundational literacy skill that is crucial for human communication and intellectual development. Bazerman (2008) highlights its indispensable role in enabling effective communication, and (Powell, 2012) (p. 12) even asserts that writing is “the most important technology in the history of the human species, except how to make fire.” Beyond its communicative function, writing is essential for fostering critical thinking, supporting learning across disciplines, and enabling individuals to organize and articulate complex ideas (Deane et al., 2008; Adler-Kassner and Wardle, 2022; Weigle, 2002). Given its significance, writing assessment plays a central role in evaluating language proficiency, communication skills, and analytical thinking. Standardized tests, such as the TOEFL (Test of English as a Foreign Language), IELTS (International English Language Testing System), and the GRE (Graduate Record Examination), include writing tasks designed to assess a test-taker’s ability to construct coherent arguments, employ appropriate vocabulary, and demonstrate grammatical accuracy.
The rapid advancement of large language models (LLMs) has significantly transformed the landscape of writing by enabling the generation of high-quality and coherent essays (OpenAI, 2024). As a result, AI-assisted writing has become increasingly prevalent in both educational and professional contexts. This widespread adoption of AI tools is reshaping how writing is created and assessed, with implications that span both the short and long term (Hao, von Davier, et al., 2024). In the short term, AI-generated essays raise serious concerns about academic integrity and test security. Students and test-takers may submit AI-assisted responses as their own work, challenging the ability of educators and testing organizations to evaluate writing skills and ensure fairness accurately. In the long term, the pervasive use of AI assistance could fundamentally alter human writing practices. Over-reliance on AI tools risks diminishing writing proficiency and critical thinking skills, as individuals may bypass the cognitive processes essential for effective communication and critical thinking. No doubt, these new developments carry significant implications for writing assessment.
In this paper, we compare the characteristics and quality of AI-generated essays, examining two key implications for writing assessment: automated essay scoring and the detection of AI-generated responses for academic integrity. The study utilizes a large set of essays generated by various LLMs in response to prompts from the high-stakes GRE Analytical Writing Assessment (https://www.ets.org/gre/test-takers/general-test/prepare/content/analytical-writing.html) developed by Educational Testing Service (ETS) and leverage the official GRE scoring rubrics implemented by both well-trained human raters and the e-rater® automated scoring engine (Attali and Burstein, 2006). We specifically aim to address the following three research questions:
  • RQ1. How do essays generated by different large language models (LLMs) differ in their characteristics and quality as compared to human-written essays?
  • RQ2. What are the limitations of current automated scoring engines when evaluating AI-generated essays compared to human raters, and how can these limitations be addressed?
@@
  • RQ3. How effective are detectors in identifying AI-generated essays using language and perplexity features when trained on essays produced by the same or different LLMs as those in the evaluation dataset?
To address RQ1, we begin by examining the linguistic and statistical characteristics of essays produced by a diverse set of high-performing LLMs as well as a random sample of human-written essays. This analysis includes evaluating adherence to prompt-specified length, language features, textual similarity, and perplexity [^1] Perplexity is a measure of how well a language model predicts a sequence of words (tokens). Mathematically, it is the exponentiated average negative log-likelihood of the predicted tokens. Lower perplexity indicates that the model assigns a higher probability to the text, suggesting the text is more likely to have been generated by the language model.. By comparing these characteristics across different models, we aim to uncover patterns that influence assessment outcomes, providing a nuanced understanding of how AI-generated texts differ from human-written responses. In response to RQ2, we assess the performance of a well-established automated essay scoring engine, e-rater®, by comparing its scoring of AI-generated essays against that of human raters. This comparison allows us to identify specific limitations of the automated scoring system, from which we propose targeted improvements to better handle AI-generated or influenced essays. Finally, to explore RQ3, we investigate the effectiveness of detectors designed to identify AI-generated essays using language and perplexity features when training data and testing data are essays generated by the same or different LLMs. Collectively, these insights advance our understanding of the implications of LLMs for writing assessments and offer practical strategies for addressing these emerging challenges.
2 Literature Review
2.1 Advance in LLMs
A language model in Natural Language Processing (NLP) refers to a statistical model that outputs a probability for a given sequence of texts (Jurafsky & Martin, 2008). Over the past few years, with the advancement of GPU-based computing power, deep neural network-based modeling techniques, and the vast amount of data available on the internet, language models have grown significantly in size, leading to the emergence of so-called large language models. LLMs are characterized by their massive parameter counts, often ranging from billions to trillions, and are trained on extensive corpora containing diverse text sources, including books, websites, and publicly available research papers (Brown et al., 2020; Kaplan et al., 2020).
Since the debut of ChatGPT, a variety of powerful LLMs have emerged, both open-source and proprietary. As of early 2025, the leading proprietary models include the GPT family from OpenAI (OpenAI, n.d.), the Gemini family from Google DeepMind (Google DeepMind, n.d.), the Grok family from xAI (xAI, n.d.), and the Claude family from Anthropic (Anthropic, n.d.). On the open-source front, several notable models have been introduced, including the Llama family from Meta (Meta AI, n.d.), the DeepSeek family (DeepSeek, n.d.), the Mistral family (Mistral, n.d.), the Qwen family from Alibaba (Alibaba Cloud, n.d.), and the Phi family from Microsoft (Microsoft, n.d.). For a comparison of their performance across tasks, readers may consult the LLM leader board (LLM Stats, n.d.). These models represent a diverse landscape of LLM capabilities, each contributing uniquely to advancements in natural language processing and AI-driven applications.
Among all the capabilities, a relevant one to this study is generating coherent and high-quality essays based on the user’s prompt. Currently, LLMs generate text using an autoregressive approach, which involves predicting and selecting each subsequent word or token sequentially based on the probability distributions learned from vast amounts of pretraining data (C.-H. Chiang and Lee, 2023; OpenAI, 2024; Ouyang et al., 2022; Touvron et al., 2023). This process begins with a prompt, which serves as the initial context, and the LLMs then generate one word at a time by evaluating the likelihood of all possible next words conditioned on the previous context, selecting the one with the highest probability. This step-by-step prediction continues iteratively, allowing the model to construct coherent and contextually appropriate sentences and paragraphs. Advanced techniques such as attention mechanisms and transformer architectures (Vaswani et al., 2017) further enhance LLMs’ ability to capture long-range dependencies between words, ensuring that the generated text remains relevant and logically consistent with the initial prompt.
2.2 Automated Scoring
Automated scoring, which involves the use of computer algorithms to automatically evaluate the constructed responses, is a critical technology used in writing assessments. The initial conception of automated scoring dates back to the 1960s (Page, 1966) and the first automated scoring system was successfully implemented in the 1990s (Page & Petersen, 1995). Several automated scoring engines have been used in operations for scoring short-constructed responses (e.g., Braun et al., 2006; Heilman & Madnani, 2015), long essays (e.g., Attali and Burstein, 2006), and and mathematical equations (Fife, 2017). Recently, the transformer architecture underlying most LLMs has shown excellent performance in automated scoring of short-constructed responses, essays, and speech ((Lottridge et al., 2022a; Lottridge et al., 2022b; Lottridge et al., 2022c).).
Automated scoring models can be broadly categorized into two main types: prompt-specific vs. generic models (Flor & Hao, 2021). Each type serves distinct purposes and is optimized for different assessment contexts, providing valuable tools for evaluating written responses efficiently and accurately. Prompt-specific models are trained on responses to a particular prompt or a closely related set of prompts. These models are fine-tuned to capture the specific expectations, vocabulary, and content relevant to a given prompt. By focusing on prompt-specific features (such as n-grams or semantic embeddings), they offer more accurate and context-sensitive scoring, ensuring that responses are evaluated against the precise criteria expected for that task. In practice, prompt-specific models are typically employed for scoring short constructed responses to open-ended questions.
In contrast, generic models are designed to operate across a wide range of topics and prompts. Unlike prompt-specific models, they focus on assessing broader constructs that are not tied to particular contexts, such as writing competency, analytical reasoning, or critical thinking. Generic models typically rely on non-contextualized features, such as grammar, organization, and logical structure, rather than features confined to specific contexts. As a result, they can effectively evaluate responses to new or unfamiliar prompts, offering greater flexibility across assessment settings. In practice, generic models are widely used for scoring longer essays in standardized testing. For example, e-rater® is a generic automated scoring model that has been used to score essays in assessments such as TOEFL and GRE. In this study, we focus on the generic model. Unless otherwise specified, all subsequent references to automated scoring refer to this type of model.
Before the advent of LLMs, automated scoring systems typically relied on a two-step process: first, extracting numerical features from texts using natural language processing (NLP) techniques, and second, mapping these features to scores through statistical or machine learning models (Attali and Burstein, 2006; J. Chen et al., 2016). However, since the introduction of LLMs, two new approaches have emerged. The first approach is end-to-end, where texts and their corresponding scores are used to fine-tune LLMs (such as BERT), establishing a direct mapping between texts and scores without the need for explicit feature extraction. The second approach involves directly instructing LLMs to score texts by providing them with scoring rubrics and a few examples (Hao, Cui, et al., 2024), which heavily relies on the in-context (few-shot) learning capabilities of LLMs, enabling them to generate scores directly based on the provided guidelines. While LLM-based approaches often outperform traditional methods in terms of accuracy, they also introduce new challenges. Specifically, the end-to-end and rubric-based approaches significantly reduce the explainability of the models, making it more difficult to understand the basis for the assigned scores (Casabianca et al., 2025). Additionally, these methods can amplify biases present in the training data, leading to bias across demographic groups (Johnson & Zhang, 2024).
2.3 Detection of AI-generated Essays
The detection of machine-generated essays initially focused on identifying texts specifically designed to deceive automated scoring systems. For instance, the Babel essay generation system (BABEL Generator, 2014) was created to generate essays intended to mislead these scoring engines. This tool accepts a list of keywords as input and uses them to construct essays that exhibit well-formed syntactic structures and advanced vocabulary. Although these essays lack genuine semantic depth, they appear linguistically complex and coherent, making it difficult for automated systems to recognize them as artificially generated. In response, Cahill et al. (2018) proposed an advisory flagging system capable of detecting such machine-generated essays by identifying patterns indicative of artificial construction.
With the advent of LLMs, the quality of generated texts has improved significantly, prompting the development of more advanced detection systems. For instance, Yan et al. (2023) demonstrated that essays generated by GPT-3 in response to large-scale writing tests could be detected with high accuracy using two distinct approaches: a language feature-based supervised learning method and an end-to-end classifier built on a fine-tuned LLM, RoBERTa (Liu et al., 2019). Following the release of ChatGPT in November 2022, detecting AI-generated texts quickly became a focal point of (Crothers et al., 2023; Hao & Fauss, 2024; Tian, 2023), and and several detectors have been developed to identify ChatGPT-generated essays in large-scale writing assessments (Hao & Fauss, 2024; Jiang, Hao, et al., 2024a, b). These detectors leverage a combination of language and perplexity-based features, achieving detection accuracies exceeding 98%. Additionally, detectors based on keystroke dynamics have also been introduced, proving effective at distinguishing authentic draft writing and copywriting by analyzing typing patterns and temporal characteristics (Deane et al., 2025; Jiang, Zhang, et al., 2024). A lingering question, however, is whether models trained on data from one LLM can reliably detect essays generated by other LLMs, especially as new models continue to emerge every few months.
@@ 3 Methods
3.1 GRE Analytical Writing and E-rater® Automated Scoring Engine
The GRE analytical writing is well known for assessing test taker’s ability to think critically and communicate complex ideas effectively. It consists of a 30-minute writing task,” analyze an issue,” which requires examinees to ponder a topic, evaluate its intricacies, and formulate a well-reasoned argument buttressed by examples and explanations. The performance is scored into six levels: fundamentally deficient (1), seriously flawed (2), limited (3), adequate (4), strong (5), and outstanding (6). More details of the scoring framework can be found in Appendix A. Overall, the assessment aims to evaluate a candidate’s ability to articulate and substantiate complex thoughts, construct and evaluate arguments, and maintain a coherent and focused discourse throughout the essay (ETS, n.d.-a).
Each essay in the GRE analytical writing is scored by a well-trained human rater and by ETS’s automated scoring engine, e-rater®. Human raters undergo a robust training and certification process: they must meet strict eligibility criteria, complete training on the scoring rubric, and pass a calibration test at the start of each scoring session to ensure scoring accuracy. During scoring, pre-scored “validation” essays are interspersed to monitor accuracy in real time; if a rater fails to maintain consistency even after retraining, they are dismissed Wendler et al. (2019). The e-rater® takes in an essay and outputs a score (in the range of 1 to 6), which is a weighted sum of a set of language features, such as grammar, usage, mechanics, styles, organization, development, and word complexity (J. Chen et al., 2016). Table 1 gives a summary of a subset of language features we focused on in this paper. These features are computed using NLP techniques, and their weights in the final score are determined by linear regression or machine learning models using a large corpus of essays scored by human raters. As such, the final e-rater® score depends on both the selection of language features and their corresponding weights based on human-scored essays. Extensive studies and literature documents that the e-rater score is well agreed with human scores Educational Testing Service (n.d.-b, n.d.-a); Williamson et al. (2012).
@@ 3.1 GRE Analytical Writing and E-rater® Automated Scoring Engine
3.2 LLM and AI Essay Generation
Since the release of ChatGPT in November 2022, LLMs have undergone an unprecedented period of rapid advancement, with increasingly powerful models emerging every few months. This study focuses on a range of LLMs, both proprietary and open-source, available between February 2023 and May 2024, a period that captures the peak of this rapid development. After this point, progress in LLMs has shifted from scaling data and model size toward enhancing reasoning capabilities, primarily through more sophisticated internal prompting techniques rather than additional training on new datasets. Within the proprietary sector, we chose the top ones, including GPT-4o from OpenAI and Gemini1.5 from Google. These closed-source models represent the forefront of AI research in the industry, offering a glimpse into the state-of-the-art capabilities in language generation. In addition to these new models, we also included some older ones, such as GPT-4 and GPT-3.5 Turbo from OpenAI and Bard from Google, which were introduced in the middle of 2023.
On the other hand, we also included some leading open-source LLMs in our evaluation. Specifically, we considered the recent Llama-3 8B model, which has been the subject of extensive research since its release (Dubey et al., 2024). In addition, we experimented with some older models from the middle of 2023, including the Llama-1 13B (Touvron et al., 2023), as well as its two finetuned variants: Vicuna (W.-L. Chiang et al., 2023) and Koala (Geng et al., 2023), both of which underwent specialized optimization processes to further refine their language generation mechanisms. The 13-billion parameter version was selected for each of these earlier models. In addition to the Llama family of models from Meta, we included another popular LLM, Mistral-7B-Instruct-v0.1, from Mistral AI, which is comparable in size to the Llama-3 model. The detailed version information of these models is presented in Appendix B.
All writing prompts of the GRE writing assessment are publicly available (ETS, n.d.-b). To alleviate the risk of biases introduced by an individual writing prompt, we randomly selected two writing prompts for this study. To prompt LLMs to generate the essays, we sent the original GRE writing prompt plus an instruction on length: “Please keep the response to about 500 words.”, where the 500 words is the typical length of GRE essays. The resulting prompts are listed below.
We generated 100 essays for each prompt from each LLM, leading to a total of 2,000 essays from the ten models. Each LLM comes with a set of hyperparameters that can be adjusted to control the text generation. Among these hyperparameters, temperature controls the randomness of token selection by scaling the logits before applying the SoftMax function. Given logits $z_{i}$, the probability of selecting token i with temperature T is $p_{i}=\frac{\exp(z_{i}/T)}{\sum_{j}\exp(z_{j}/T)}$. Lower temperatures make the output more deterministic by sharpening the distribution while higher temperatures increase the randomness, leading to more diverse but potentially less coherent texts.
For comparison purposes, we set the temperature to 0.4 (the default value used in earlier versions of the web-based ChatGPT to strike a balance between diversity and quality) for each model, except for Bard, where this parameter configuration was not accessible. Here, we caution that different LLMs may use varying scaling for their temperature, meaning that the same numerical value may affect the randomness and diversity of responses differently across models. In addition to AI-generated essays, we randomly sampled 100 human-written essays for each of the two prompts from the 2021 GRE writing test, a period before ChatGPT-like AI systems were available. This ensures that the selected essays were authentically human-written. We chose 100 human essays to align with the number of AI-generated essays per prompt per LLM, facilitating direct comparison. The decision to generate 100 AI essays per prompt per LLM was made to strike a balance between capturing sufficient variability in performance and maintaining feasibility for analysis. Together, these human- and AI-written essays provide a comparable and well-controlled basis for evaluation as well as training data for developing detection models.
3.3 Essay Characteristics
To address RQ1 (how essays generated by different large language models (LLMs) differ in their characteristics and quality compared to human-written essays), we examine a set of selected features of the AI-generated essays. These characteristics were chosen because they are highly relevant to the other two research questions, e.g., automated scoring and the detection of AI-generated essays.
Essay Length. We compute the length of the generated essay to check how well each LLM adheres to the 500 words in the prompt.
Essay Similarity. We compute the pairwise similarity between essays generated by the same LLM to assess the variation in the LLM’s outputs under the same prompt. There are two commonly known types of similarities between texts: verbatim similarity, which describes how closely two texts match in exact wording or surface structure, and semantic similarity, which shows how similar two texts are in meaning, regardless of specific word choices. Comparing these two types of similarities for pairwise essays generated by LLMs under the same prompt could reveal how well the LLMs maintain meaning across different expressions of the same content, having important implications for automated scoring and the detection of AI-generated essays.
For semantic similarity, we first convert each essay into a document vector using a popular neural embedding model jina-embedding-v3 model (Sturua et al., 2024). This embedding model was chosen based on its high performance on text similarity benchmarks of the MTEB English leaderboard and its superior ability to handle long documents. Once the essays are converted to vectors, we compute the cosine similarity between pairs of vectors corresponding to essays generated from the same LLM under the same prompt. Cosine similarity measures the similarity between two vectors by comparing the angle between them, regardless of their magnitude. It ranges from –1 to 1, where values close to 1 indicate strong similarity (vectors pointing in the same direction), values near 0 suggest little to no similarity (orthogonal vectors), and values near –1 reflect strong dissimilarity or opposite direction. In practice, for text embeddings, cosine similarity above 0.8 typically indicates high similarity, around 0.5 suggests moderate similarity, and below 0.2 indicates minimal similarity.
For verbatim similarity, we represent each essay as a trigram vector, which encodes the frequencies of overlapping three-word sequences in the text. This representation is effective in capturing textual similarities for test security applications (Choi et al., 2024). Based on the trigram vectors, we computed pairwise cosine similarities for each pair of essays generated by the same LLM for the same prompt.
Language Features. We compute the language features as shown in Table 1 for each essay using e-rater® and then compare the statistics of these features across the essays generated by different LLMs.
Perplexity Distribution. Perplexity characterizes the likelihood of a sequence of words (in an essay) being generated from a given language model and plays a crucial role in detecting AI-generated texts. Most LLMs used in this study do not allow the direct computation of perplexity. Therefore, we selected GPT-2 as a common language model to calculate the perplexity of essays generated by all LLMs. As such, the perplexity here is a relative measure of how likely the essays have been generated by GPT-2. Prior research has demonstrated that perplexity, computed based on a common language model, is highly informative for detecting AI-generated essays. (e.g. Hao & Fauss, 2024, Jiang, Hao, et al., 2024a).
3.4 Essay Scores
We recruited expert human raters from the GRE program to score the 200 essays (100 from each of the two writing prompts) generated by each LLM and humans. The raters applied the same scoring rubric and procedures used in the operational GRE program, assigning each essay a score on a scale from 1 to 6 (see Appendix A for score definitions). The source of the essay, i.e., written by humans or AI, has been deliberately hidden from the human raters, so they were unaware the essays were generated by AI. In addition to human scores, we used the e-rater® engine to score each AI-generated essay and compared its scores with those assigned by human raters. The discrepancies between the two sets of scores highlight the limitations of the current automated scoring system in evaluating AI-generated content. They also highlight opportunities for improvement, helping to better prepare automated scoring engines for the growing prevalence of AI-generated and AI-assisted texts, which is exactly the focus of RQ2.
3.5 Detection of AI-generated Essays
As the quality of essays generated by LLMs continues to improve, concerns are rising in education about students submitting AI-generated assignments, which poses a significant challenge to academic integrity. These concerns have become a significant risk, even in high-stakes testing, which began with the rise of remotely proctored exams following the COVID-19 pandemic (Hao & Fauss, 2024). The detection of AI-generated essays in large-scale assessments has been a widely discussed topic Hao & Fauss (2024); Jiang, Hao, et al. (2024a); Yan et al. (2023), with most previous work focusing on detecting essays from a single LLM. However, with new and more powerful LLMs emerging every few months, it becomes increasingly difficult to identify which LLMs are being used. This highlights the importance of examining the detectability of AI-generated essays across multiple LLMs, which is the RQ 3 of this study.
Specifically, we consider two scenarios: within-model detectability and cross-model detectability. In the within-model scenario, we assess how well classifiers trained on essays generated by a specific LLM can detect essays produced by that same LLM. In the cross-model scenario, we examine how effectively a classifier trained on data from one LLM can detect essays generated by other LLMs. In both cases, the classifiers were developed using the seven language features in Table 3 and perplexity-based features (including the whole essay perplexity score, the mean, median, maximum, minimum, and 10, 20, 30, …, 90 percentiles of the sentence perplexity in each essay), respectively. We experimented with multiple popular classifiers, including XGBoost (T. Chen and Guestrin, 2016), Random Forest (Breiman, 2001), Gradient Boosted Machine (Friedman, 2001) and Support Vector Machine (Cortes & Vapnik, 1995) . We found that XGBoost consistently yielded the best results. So, we only show the results of the XGBoost classifier.